An electric field of strength E 4700 N/C is directed along the +x-axis as shown. What is the magnitude of the potential difference between point B and point A if the distance between these two points is 10 cm? Submit Answer Tries 0/2 An electron is initially at rest at point B. How much kinetic energy (in electron-Volts, or ev) will the electron gain by moving to the higher potential at point A? eV Submit Answer Tries 0/2 What is the kinetic energy of the electron at point A measured in Joules? Submit Answer Tries 0/2
Answer #1

a)

Relation between electric field and potential difference is given as,

E = \left | \frac{dV}{dx} \right |

So the potential difference between the points A and B is,

dV = E \times dx

dV = 4700 \times 0.1

dV = 470 volt

b) Change in energy is given by,

dw= q \times dV

dw= e \times 470V = 470 eV

c) Kinetic energy in Joules

dw=1.602\times 10^{-19}\times 470 = 7.53\times 10^{-17}J